OpenStudy (anonymous):

Differentiate f(x)=-5(12^x)

OpenStudy (anonymous):

\[f(x)=b^x\] \[f'(x)=b^x\times \ln(b)\]

OpenStudy (lgbasallote):

Proof: \(\large \text{let y} = b^x\) \(\ln y = x \ln b\) by implicit differentiation \(\Large \frac{y'}{y} = \ln b\) \(y' = y \lnb\) \(\Large y' = b^x \ln b\)

OpenStudy (lgbasallote):

that second to last line is \(\large y' = y \ln b\)